Explain in complete sentenceswhy these two equqtions are equivalent.
A)x + 3 = 9

B)x - 3 = 3
Thanks for the help

Found 2 solutions by psbhowmick, JoelSchwartz:
Answer by psbhowmick(878)

A) x + 3 = 9.
Subtract 6 from both sides.
Thus, x + 3 - 6 = 9 - 6
or x + (3 - 6) = 3 [collecting similar terms together; 3 and 6 are similar, can be added or subtracted]
or x - 3 = 3

B) x - 3 = 3

Thus these two equations are same or equivalent.

Answer by JoelSchwartz(130)

Both equations have the same x value.
x+3-9=9-9
x-6=0
x-3=3-3
x-3-3=0
x-6=0
x-6=x-6
x-6+6=x-6+6
x=x